# Knowledge Adaptation as Posterior Correction

### v1

@@ 4.3 Posterior Correction for Continual Learning
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})&=\mathcal{L}[y_{i},\hat{y}_{i}(\boldsymbol{\theta})]-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla\mathcal{L}[y_{i},\hat{y}_{i}]\right]\\ &=-y_{i}f_{i}(\boldsymbol{\theta})+A(\boldsymbol{\theta})-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla f_{i}(\boldsymbol{\theta})\left(\hat{y}_{i}(\boldsymbol{\theta})-y_{i}\right)\right]\\ &\approx-\hat{y}_{i|t}f_{i}(\boldsymbol{\theta})+A(\boldsymbol{\theta})+r_{i|t}f_{i}(\boldsymbol{\theta})-r_{i|t}\boldsymbol{\theta}^{\top}\nabla f_{i}(\mathbf{m}_{t})\\ &=\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]+r_{i|t}\left[f_{i}(\boldsymbol{\theta})-\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})\right]+\text{const.}\end{split}$ (31)
The first and second line follows from the definition of the loss and site. In the third line, we subtract and add $\hat{y}_{i|t}f_{i}(\boldsymbol{\theta})$, then rearrange to write it in terms of the residual $r_{i|t}=\hat{y}_{i|t}-y_{i}$. The approximation is due to the $1^{\text{st}}$-order delta method to simply the expectation with respect to $q_{t}$ over $f_{i}(\boldsymbol{\theta})$ in the last term. The final line is obtained by using $\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})=(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla f_{i}(\mathbf{m}_{t})$ with a term that is constant with respect to $\boldsymbol{\theta}$. The only assumption made is the delta method for the expectation which might be useful for Bayesian cases but otherwise can be safely ignored. Other than that, the expression holds for general models, such as, neural networks.
@@
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{full}}_{i|t}(\boldsymbol{\theta})&=\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})-\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\mathbf{H}_{i|t}(\boldsymbol{\theta}-\mathbf{m})\\ &\approx\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]+r_{i|t}\left[f_{i}(\boldsymbol{\theta})-\hat{f}^{\text{quad}}_{i}(\boldsymbol{\theta})\right]-\mbox{$\frac{1}{2}$}\beta_{i|t}\|\nabla f_{i}^{\text{lin}}\|^{2},\end{split}$ (32)
where we denote $\hat{f}_{i}^{\text{quad}}(\boldsymbol{\theta})=\hat{f}_{i}^{\text{lin}}(\boldsymbol{\theta})+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla^{2}f_{i}(\mathbf{m}_{t})(\boldsymbol{\theta}-\mathbf{m}_{t})$, and $\beta_{i|t}=\mathbb{E}_{q_{t}}[A^{\prime\prime}(f_{i}(\boldsymbol{\theta}))]$. A derivation is in App. A which involves writing the Hessian of the loss in terms of the Generalized Gauss-Newton (GGN) matrix and the Hessian of the model output. Applying the delta method, then gives this approximation. The linearization error is now reduced due to the use of a better quadratic surrogate, as opposed to the linear surrogate used in Eq. 31. The last term further reduces the error and is due to the GGN matrix.
@@
The second term essentially attempts to improve the current posterior $q_{t}$ by removing the stale $\beta_{i|t}$ and replace it by the fresh ones by using the prediction mismatch term. By using a more flexible posterior, we reduce the correction required for accurate adaptation. This idea was first used in K-priors by Khan and Swaroop (2021, Fig. 3a) through gradient matching but it is also connected to PVI where we replace old surrogates by new ones during federated learning. In general, all such procedures are generalized via posterior correction. We next show that the correction term generalizes K-priors to exponential-family posteriors.
@@ 4.4 Knowledge-Adaptation Priors as Posterior Correction
The prior shown in Eq. 6 is a generalization of the K-prior by Khan and Swaroop (2021). We will now show this for the K-prior presented in Khan and Swaroop (2021, Eq. 8) for a linear model $f_{i}(\boldsymbol{\theta})=\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}$ over the Bregman loss $\ell_{i}(\boldsymbol{\theta})=\mathcal{L}(y_{i},\hat{y}_{i}(\boldsymbol{\theta}))$ and quadratic regularizer $\ell_{0}(\boldsymbol{\theta})=\mbox{$\frac{1}{2}$}\delta\|\boldsymbol{\theta}\|^{2}$. We have already derived the correction terms for $i=1,2,\ldots,t$. Therefore, we just need the correction term for $\ell_{0}$ and the KL term. We derive this below.
@@
$\begin{split}\ell_{0}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{0|t}(\boldsymbol{\theta})&=\ell_{0}(\boldsymbol{\theta})-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{0}+\nabla\log h]+\log h(\boldsymbol{\theta})\\ &=\mbox{$\frac{1}{2}$}\delta\|\boldsymbol{\theta}\|^{2}-\delta\boldsymbol{\theta}^{\top}\mathbf{m}_{t}+\boldsymbol{\theta}^{\top}\mathbf{m}_{t}-\mbox{$\frac{1}{2}$}\|\boldsymbol{\theta}\|^{2}\\ &=\mbox{$\frac{1}{2}$}\delta\|\boldsymbol{\theta}-\mathbf{m}_{t}\|^{2}-\mbox{$\frac{1}{2}$}\|\boldsymbol{\theta}-\mathbf{m}_{t}\|^{2}+\text{const.}\\ \end{split}$ (34)
@@
$\mathbb{E}_{q}[\ell_{0}-\hat{\ell}_{0|t}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]=\mbox{$\frac{1}{2}$}\delta\|\mathbf{m}-\mathbf{m}_{t}\|^{2}+\text{const.}$. (35)
@@
$\mathcal{K}_{t}^{\text{iso}}(\mathbf{m})=\sum_{i=t}^{t}\mathbb{E}_{q}\left(\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]\right)+\delta\mbox{$\frac{1}{2}$}\|\mathbf{m}-\mathbf{m}_{t}\|^{2}$ (36)
which is exactly Khan and Swaroop (2021, Eq. 8) with respect to $\mathbf{m}$, if we use the delta method: $\hat{y}_{i|t}=\mathbb{E}_{q_{t}}[\hat{y}_{i}(\boldsymbol{\theta})]\approx\hat{y}_{i}(\mathbf{m}_{t})$ and $\mathbb{E}_{q}\left(\mathcal{L}\left[\hat{y}_{i}(\mathbf{m}_{t}),\,\hat{y}_{i}(\boldsymbol{\theta})\right]\right)\approx\mathcal{L}[\hat{y}_{i}(\mathbf{m}_{t}),\,\hat{y}_{i}(\mathbf{m})]$.
@@
The two sets have to be disjoint so as to not double count the contribution of prediction matching. The benefit of using the second term above is that those examples do not need labels and they can be summarized using arbitrary input locations, for instance, similar to a core-set or set of inducing inputs. Daxberger et al. (2023) used a memory construction similarly to the above for K-prior to get good improvement on the ImageNet dataset. Similarly, Pan et al. (2020) and Khan and Swaroop (2021) used a quantity similar to $\beta_{i|t}$ to pick examples to do prediction matching on. They all show consistent improvement and point this as a viable option to design better memory replay methods.
We now briefly discuss connections to other approaches covered under the K-prior framework. For example Knowledge Distillation (KD) (Hinton et al., 2015) considers a teacher-student learning scenario which can also be written as a posterior correction. The KD objective uses a convex combination for $\gamma\in(0,1)$ which can be rewritten as,
@@
Posterior correction is also closely related to dual approaches in SVM, such as, incremental Support Vector Machines (Cauwenberghs and Poggio, 2001; Liang and Li, 2009), Similarity Control (Vapnik and Izmailov, 2015), and their extensions to neural networks, for example, (Lopez-Paz and Ranzato, 2017). The correction term can be written in a trust-region form with constraints $\mathbb{E}_{q}[\ell_{i}]=\mathbb{E}_{q}[\hat{\ell}_{i}]$. The dual problem can be connected to posterior correction via a Lagrangian. The continual learning case can be written in a similar fashion.
@@ 4.5 Posterior Correction for Sequential and Online Variational Inference
It is well-known that the recursive elegance of Bayes’ rule is lost when approximations are used. The problem of an incorrect $\hat{q}_{t+1}$ when Eq. 2 is used occurs in almost all approximate inference problems. This is true for even the simplest variational (Bayesian) learning on conjugate models where mean-field approximations are used (Sato, 2001; Ghahramani and Attias, 2000; Honkela and Valpola, 2003; Hoffman et al., 2010; Wang et al., 2011), but also more recent variants of those Broderick et al. (2013); Nguyen et al. (2018); Zeno et al. (2018); Chérief-Abdellatif et al. (2019); Jones et al. (2024). Other approximations also suffer from this issue, for example, those proposed in Opper and Winther (1999); Winther and Solla (1998); Heskes and Zoeter (2002); Csató and Opper (2002), among many others. We expect similar problems to exist even when amortized inference is used (Archer et al., 2015; Kim et al., 2020; Krishnan et al., 2017; Campbell et al., 2021).
The posterior correction is a useful approach to fix such issues in sequential and online variational inference. As an example, we show that an existing method of Bui et al. (2017) is in fact an example of posterior correction for Sparse Variational Gaussian Process (SVGP). The method attempts to handle a challenging case of updating both the inducing input as well as hyperparameters but we will consider a simpler version to show similarities to posterior correction. SVGP uses a set of inducing variables, denoted by $\mathbf{u}$, to model the function $f$. To keep things simple, we will use a notation that fits in our framework. We will denote the posterior by $q(\mbox{$\mbox{$\mathbf{u}$}$},f)$ and the prior by $p_{0}^{\delta}(\mbox{$\mbox{$\mathbf{u}$}$},f)$ where $\delta$ is the set of kernel hyperparameter. We will denote the negative likelihoods by $\ell_{t}(f)$ which are assumed to be Gaussian. With this notation, we assume that $q_{t}$ and $\delta_{t}$ is given and our goal is to update $q_{t+1}$ as well as $\delta_{t+1}$, but we assume the inducing variables $\mathbf{u}$ remain at the same locations.
Because the model is conjugate, the sites are not approximate, that is, ^$\hat{\ell}_{i|t}=\ell_{i}$. With this, we can write the following (denoting the normalizing constant by $\mathcal{Z}_{t}(\delta_{t})$),
@@
$\begin{split}q_{t+1}&=\arg\min_{q\in\mathcal{Q}}\,\,\sum_{i=1}^{t+1}\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]+\log\mathcal{Z}_{t+1}(\delta)\\ &=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\mathbb{E}_{q}[\ell_{0}^{\delta}-\ell_{0}^{\delta_{t}}]+\log\frac{\mathcal{Z}_{t+1}(\delta)}{\mathcal{Z}_{t}(\delta_{t})}\\ &=\arg\min_{q\in\mathcal{Q}}\,\,\log\frac{\mathcal{Z}_{t+1}(\delta)}{\mathcal{Z}_{t}(\delta_{t})}+\mathbb{E}_{q}\left[\log\left(\frac{p_{0}^{\delta_{t}}}{p_{0}^{\delta}}\frac{q}{q_{t}}\frac{1}{e^{-\ell_{t+1}}}\right)\right]\end{split}$ (37)
where in the first line we added the explicit normalizing constant to show dependency on $\delta$. The last line is written to show equivalence to Bui et al. (2017, Eq. 5, $2^{\text{nd}}$ line). Modification to handle non-conjugate likelihood is straightforwardly obtained by adding back the corrections for all the past $\ell_{i}$. The procedure can also benefit from adding memory replay, similarly to previous sections. A recent extension by Chang et al. (2023) explores such directions using the dual representations and finds good improvements. Additionally, a procedure by Adam et al. (2021) is also a special instance of posterior correction to speed up hyperparameter learning in GPs.

### v3

@@ 4.3 Posterior Correction for Continual Learning
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})&=\mathcal{L}[y_{i},\hat{y}_{i}(\boldsymbol{\theta})]-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla\mathcal{L}[y_{i},\hat{y}_{i}(\boldsymbol{\theta})]\right]\\ &=-y_{i}f_{i}(\boldsymbol{\theta})+A(f_{i}(\boldsymbol{\theta}))-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla f_{i}(\boldsymbol{\theta})\left(\hat{y}_{i}(\boldsymbol{\theta})-y_{i}\right)\right]\\ &=-\hat{y}_{i|t}f_{i}(\boldsymbol{\theta})+(\hat{y}_{i|t}-y_{i})f_{i}(\boldsymbol{\theta})+A(f_{i}(\boldsymbol{\theta}))-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla f_{i}(\boldsymbol{\theta})\left(\hat{y}_{i}(\boldsymbol{\theta})-y_{i}\right)\right]\\ &=-\hat{y}_{i|t}f_{i}(\boldsymbol{\theta})+r_{i|t}f_{i}(\boldsymbol{\theta})+A(f_{i}(\boldsymbol{\theta}))-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla f_{i}(\boldsymbol{\theta})\left(\hat{y}_{i}(\boldsymbol{\theta})-y_{i}\right)\right]\\ &\approx-\hat{y}_{i|t}f_{i}(\boldsymbol{\theta})+r_{i|t}f_{i}(\boldsymbol{\theta})+A(f_{i}(\boldsymbol{\theta}))-\boldsymbol{\theta}^{\top}{\nabla f_{i}(\mathbf{m}_{t})\left(\mathbb{E}_{q_{t}}[\hat{y}_{i}(\boldsymbol{\theta})]-y_{i}\right)}\\ &=-\hat{y}_{i|t}f_{i}(\boldsymbol{\theta})+A(f_{i}(\boldsymbol{\theta}))+r_{i|t}f_{i}(\boldsymbol{\theta})-r_{i|t}\boldsymbol{\theta}^{\top}\nabla f_{i}(\mathbf{m}_{t})\\ &=\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]+r_{i|t}[f_{i}(\boldsymbol{\theta})-\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})]+\text{const.}\end{split}$ (44)
The first and second line follows from the definition of the loss and site. In the third line, we subtract and add $\hat{y}_{i|t}f_{i}(\boldsymbol{\theta})$ where $\hat{y}_{i|t}=\mathbb{E}_{q_{t}}[\hat{y}_{i}(\boldsymbol{\theta})]$. Then, we write it in terms of the residual $r_{i|t}=\hat{y}_{i|t}-y_{i}$ in the fourth line. In the fifth line, we use the first-order delta method in the last term approximating $\nabla f_{i}(\boldsymbol{\theta})\approx\nabla f_{i}(\mathbf{m}_{t})$ and move the expectation with respect to $q_{t}$ inside over to $\hat{y}_{i}(\boldsymbol{\theta})$. The last two lines are simple rearrangements and written in terms of $\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})=(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla f_{i}(\mathbf{m}_{t})$ (with a term that is constant with respect to $\boldsymbol{\theta}$). The only assumption made is the delta method for the expectation which might be useful for Bayesian cases but otherwise can be safely ignored. Other than that, the expression holds for general models, such as, neural networks.
@@
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{full}}_{i|t}(\boldsymbol{\theta})&=\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})-\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\mathbf{H}_{i|t}(\boldsymbol{\theta}-\mathbf{m}_{t})\\ &\approx\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]+r_{i|t}\left[f_{i}(\boldsymbol{\theta})-\hat{f}^{\text{quad}}_{i}(\boldsymbol{\theta})\right]-\mbox{$\frac{1}{2}$}\beta_{i|t}\|\nabla f_{i}^{\text{lin}}\|^{2},\end{split}$ (45)
where we denote $\hat{f}_{i}^{\text{quad}}(\boldsymbol{\theta})=\hat{f}_{i}^{\text{lin}}(\boldsymbol{\theta})+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla^{2}f_{i}(\mathbf{m}_{t})(\boldsymbol{\theta}-\mathbf{m}_{t})$, and $\beta_{i|t}=\mathbb{E}_{q_{t}}[A^{\prime\prime}(f_{i}(\boldsymbol{\theta}))]$. A derivation is in App. C which involves writing the Hessian of the loss in terms of the Generalized Gauss-Newton (GGN) matrix and the Hessian of the model output. Applying the delta method, then gives this approximation. The linearization error is now reduced due to the use of a better quadratic surrogate, as opposed to the linear surrogate used in Eq. 44. The last term further reduces the error and is due to the GGN matrix.
@@
The second term essentially attempts to improve the current posterior $q_{t}$ by removing the stale $\beta_{i|t}$ and replace it by the fresh ones by using the prediction mismatch term. By using a more flexible posterior, we reduce the correction required for accurate adaptation. This idea was first used in K-priors by khan2021knowledge through gradient matching but it is also connected to PVI where we replace old surrogates by new ones during federated learning. In general, all such procedures are generalized via posterior correction. We next show that the correction term generalizes K-priors to exponential-family posteriors.
@@ 4.4 Knowledge-Adaptation Priors as Posterior Correction
The prior shown in Eq. 6 is a generalization of the K-prior by khan2021knowledge. We will now show this for the K-prior presented in khan2021knowledge for a linear model $f_{i}(\boldsymbol{\theta})=\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}$ over the Bregman loss $\ell_{i}(\boldsymbol{\theta})=\mathcal{L}(y_{i},\hat{y}_{i}(\boldsymbol{\theta}))$ and quadratic regularizer $\ell_{0}(\boldsymbol{\theta})=\mbox{$\frac{1}{2}$}\delta\|\boldsymbol{\theta}\|^{2}$. We have already derived the correction terms for $i=1,2,\ldots,t$. Therefore, we just need the correction term for $\ell_{0}$ and the KL term. We derive this below.
@@
[…]
@@
$\mathbb{E}_{q}[\ell_{0}-\hat{\ell}_{0|t}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]=\mbox{$\frac{1}{2}$}\delta\|\mathbf{m}-\mathbf{m}_{t}\|^{2}+\text{const.}$ (48)
@@
$\mathcal{K}_{t}^{\text{iso}}(\mathbf{m})=\sum_{i=1}^{t}\mathbb{E}_{q}\left(\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]\right)+\delta\mbox{$\frac{1}{2}$}\|\mathbf{m}-\mathbf{m}_{t}\|^{2}$ (49)
which is exactly khan2021knowledge with respect to $\mathbf{m}$, if we use the delta method: $\hat{y}_{i|t}=\mathbb{E}_{q_{t}}[\hat{y}_{i}(\boldsymbol{\theta})]\approx\hat{y}_{i}(\mathbf{m}_{t})$ and $\mathbb{E}_{q}\left(\mathcal{L}\left[\hat{y}_{i}(\mathbf{m}_{t}),\,\hat{y}_{i}(\boldsymbol{\theta})\right]\right)\approx\mathcal{L}[\hat{y}_{i}(\mathbf{m}_{t}),\,\hat{y}_{i}(\mathbf{m})]$.
@@
The two sets have to be disjoint so as to not double count the contribution of prediction matching. The benefit of using the second term above is that those examples do not need labels and they can be summarized using arbitrary input locations, for instance, similar to a core-set or set of inducing inputs. daxberger2023improving used a memory construction similarly to the above for K-prior to get good improvement on the ImageNet data set. Similarly, pan2020continual and khan2021knowledge used a quantity similar to $\beta_{i|t}$ to pick examples to do prediction matching on. They all show consistent improvement and point this as a viable option to design better memory replay methods.
We now briefly discuss connections to other approaches covered under the K-prior framework. For example Knowledge Distillation (KD) (hinton2015distilling) considers a teacher-student learning scenario which can also be written as a posterior correction. The KD objective uses a convex combination for $\gamma\in(0,1)$ which can be rewritten as,
@@
Posterior correction is also closely related to dual approaches in SVM, such as, incremental Support Vector Machines (cauwenberghs2001incremental; liang2009incremental), Similarity Control (vapnik2015learning), and their extensions to neural networks, for example, (lopez2017gradient). The correction term can be written in a trust-region form with constraints $\mathbb{E}_{q}[\ell_{i}]=\mathbb{E}_{q}[\hat{\ell}_{i}]$. The dual problem can be connected to posterior correction via a Lagrangian. The continual learning case can be written in a similar fashion.
@@ 4.5 Posterior Correction for Sequential and Online Variational Inference
It is well-known that the recursive elegance of Bayes’ rule is lost when approximations are used. The problem of an incorrect $\hat{q}_{t+1}$ when Eq. 2 is used occurs in almost all approximate inference problems. This is true for even the simplest variational (Bayesian) learning on conjugate models where mean-field approximations are used (sato2001online; ghahramani2000online; honkela2003line; hoffman2010online; wang2011online), but also more recent variants of those broderick2013streaming; cuong2018; zeno2018task; cherief2019generalization; jones2024bayesian. Other approximations also suffer from this issue, for example, those proposed in opper1999bayesian; winther1998optimal; Heskes2002ExpectationPF; csato2002sparse, among many others. We expect similar problems to exist even when amortized inference is used (archer2015black; kim2020variational; krishnan2017structured; campbell2021online).
The posterior correction is a useful approach to fix such issues in sequential and online variational inference. As an example, we show that an existing method of bui2017streaming is in fact an example of posterior correction for Sparse Variational Gaussian Process (SVGP). The method attempts to handle a challenging case of updating both the inducing input as well as hyperparameters but we will consider a simpler version to show similarities to posterior correction. SVGP uses a set of inducing variables, denoted by $\mathbf{u}$, to model the function $f$. To keep things simple, we will use a notation that fits in our framework. We will denote the posterior by $q(\mbox{$\mbox{$\mathbf{u}$}$},f)$ and the prior by $p_{0}^{\delta}(\mbox{$\mbox{$\mathbf{u}$}$},f)$ where $\delta$ is the set of kernel hyperparameter. We will denote the negative likelihoods by $\ell_{t}(f)$ which are assumed to be Gaussian. With this notation, we assume that $q_{t}$ and $\delta_{t}$ is given and our goal is to update $q_{t+1}$ as well as $\delta_{t+1}$, but we assume the inducing variables $\mathbf{u}$ remain at the same locations.
Because the model is conjugate, the sites are not approximate, that is, $\hat{\ell}_{i|t}=\ell_{i}$. With this, we can write the following (denoting the normalizing constant by $\mathcal{Z}_{t}(\delta_{t})$),
@@
$\begin{split}q_{t+1}&=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\sum_{i=1}^{t+1}\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]+\log\mathcal{Z}_{t+1}(\delta)\\ &=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\log\frac{\mathcal{Z}_{t+1}(\delta)}{\mathcal{Z}_{t}(\delta_{t})}+\mathbb{E}_{q}\left[\log\left(\frac{p_{0}^{\delta_{t}}}{p_{0}^{\delta}}\frac{q}{q_{t}}\frac{1}{e^{-\ell_{t+1}}}\right)\right]\end{split}$ (50)
where in the first line we added the explicit normalizing constant to show dependency on $\delta$. The last line is written to show equivalence to bui2017streaming. Modification to handle non-conjugate likelihood is straightforwardly obtained by adding back the corrections for all the past $\ell_{i}$. The procedure can also benefit from adding memory replay, similarly to previous sections. A recent extension by ChangVJSK23 explores such directions using the dual representations and finds good improvements. Additionally, a procedure by adam2021dual is also a special instance of posterior correction to speed up hyperparameter learning in GPs.